substitution
a - 4b = 2
2a - 5b = 1
with steps too

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

1a+4b=2 .............1
-5b=1-2a ---------2
/-5=
b=-0.2+0.4a
Plug the value of y in (1)
1a+4(-0.2+0.4a)=2
1a-0.8+1.6a=2
1a+1.6a=2+0.8
2.6a=2.8
/2.6
a=1.08
Plug the value of x in (1)
1a+4b=2
1*1.08+4b=2
1.08+4b=2
4b=0.92
/4
b=0.23

Answer by ikleyn(53937)

.
substitution
a - 4b = 2
2a - 5b = 1
~~~~~~~~~~~~~~~~~~~~~


        The solution in the post by @mananth is  INCORRECT.

        His answer  a = 1.08,  b = 0.23  is  INCORRECT.   These values do not satisfy second equation.

        So,  I came to solve the problem in a right way.


From first equation, express  a = 4b+2 and substitute it into the second equation

    2(4b+2) - 5b = 1,

    8b + 4 - 5b = 1,

    3b = 1 - 4 = -3,

     b = -3/3 = -1.


Then  a = 4b+2 = 4*(-1) + 2 = -2.


ANSWER.  The solution is  a = -2,  b = -1.

Solved correctly.